andrew can paint the neighbor's house 3 times as fast as bailey. atahe year Andre and Bailey worked together, it took them 5 days. How long would it take each to paint the house?
Answer by richwmiller(17219)

1/a+1/b=1/x
a=b*3
1/b*3+1/b=1/5
1/b*3+3/3*b=1/5
4/b*3=1/5
3*b=20
b=6.66666667
a=6.66666667*3
a=20 b=6.66666667
check
1/a+1/b=1/x
1/20+1/6.66666667=1/5
0.05+0.15=1/5
0.2=1/5
0.2=0.2
x=5 hours working together
a=20 hours working alone
b=6.66666667 hours working alone
ok